OpenStudy (anonymous):

suppose that p/q=\[\sqrt{3}\], what is the value of \[\frac{2p}{p-q}\]? Express the answer in simplest radical form.

OpenStudy (anonymous):

you can assume q=1 and p=sqrt(3)

OpenStudy (asnaseer):

sorry - that should be:\[\frac{p-q}{2p} = \frac{p}{2p}-\frac{q}{2p}=\frac{1}{2}-\frac{1}{2\sqrt{3}}\]

OpenStudy (asnaseer):

does it make sense?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks

OpenStudy (asnaseer):

yw